Home‎ > ‎Reports‎ > ‎Crew104‎ > ‎

day07


Crew 104 Log Book for April 29, 2011

Photos of the day





Commander's Check-In Report

Jon Clarke Reporting

Crew Physical Status: Good

Time Departed/Returned from EVA: 10:30 drill serving crew left, returned 12:30.  Field sampling crew left at 3:00 pm, returned 8:00 pm

Brief Narrative of Field Mission Results:  Drill repairs abandoned because of weather (cold, wind, blowing sand). Field sampling successful collected samples for organic extraction from Summerville and Salt Wash.  Major hitch when keys were locked in hire vehicle, requiring external assistance.  This resulted in late return.

EVA Data/Interpretations:  Repair of MARTE drill still pending.  Samples collected from bedded gypsum, venin gypusm, nodular gypsum and host shales and sandstones of Summerville and from abandoned channel fill sandstones and shales and from petrified wood in Salt Wash.

Engineering/Hab Maintenance: Repair of battery terminal of Yamaha ATV in progress.

Report Transmission Schedule: Engineering and Commander's report will be submitted later. 

Plans for Tomorrow:  Further repair of MARTE, if parts arrive.  Further repair of ATV.  Ongoing lab extraction of DNA and organics.  All these are weather permitting.  The forecast is for more wind and lower temperatures (freezing alert overnight and 12 degrees tomorrow).

Inventory:  Run out of potato flakes

Support Requested: None

Miscellaneous:  Future crews that using additional hire vehicles need to be alerted to the risk of keys being locked in vehicles by self locking systems.


Commander's Report

Jon Clarke Reporting

This has been quite a day.  Having diagnosed the MARTE problem as mechanical we dispatched the drilling team of David and Jhony to work on the drill while more specialized tools to arrive.  However the strong wind, with steady speeds of nearly 40 kph and with much stronger gusts, meant that this had to be abandoned by lunch time.  Meanwhile the biology team of Julio and Luisa worked away on the extraction of organics and DNA from samples and Carol worked on XRD analysis of mineral specimens.  Sarah collected video interviews to send to the Mars 500 crew at IBMP in Moscow, now on the "homeward leg" of their ground-breaking 520 days simulated Mars mission.  One of their crew, Diego Urbina, is an MDRS veteran. Your correspondent, the humble XO and hairy-knuckled geologist, spent the morning working on reports and field notes.

After a hot lunch, doubly welcome because of the cold weather and the near sandstorm conditions outside, a sampling team was put together to collect further samples from a former quarry near Hanksville.  The team - Carol, David, Julio and myself -set off in a hire vehicle and successfully collected a range of samples from the Summerville Formation and Salt Was Member for further analysis. 

The car keys were left in the vehicle as a safety precaution against loosing them.  Unfortunately the Chevrolet Suburban's security system, HAL-like, locked the doors after us, requiring the visit of a technician from Green River.  Future crews who hire additional vehicles should be advised to keep their keys with them, regardless of vehicles types, as this is an issue with many (though not all) modern vehicles.

So a late dinner tonight and then a review of the home movies that Sarah has been preparing to send to the Mars 500 crew.

Tomorrows plans are very weather dependent.  The forecast is for subzero temperatures overnight and only 12 degrees C during the day, and more wind.

This is crew 104 signing off.



Engineering Report

David Willson Reporting

Generator/Electricity:

Kitty (Used? Yes/No): Used Continuously and not stopped

Honey (Diesel Generator) (Used? Yes/No): Not Used

Wendy (Gas Generator) (Used? Yes/No): Not Used

Battery Charge Level: Not Checked

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

Generator operating very well continuously and not stopped


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 22.5% in Tank

Propane Tank (Full/0.75/0.5/0.25/Empty): 32%

Gasoline Tank (Full/0.75/0.5/0.25/Empty): N/A

Notes/Comments:


Oil:

Kitty Oil Quantity (# of Quarts in storage): Full

Honey Oil Quantity (# of Quarts in storage): Still Very Low, Will be toped up tomorrow

ATV Oil Quantity (# of Quarts in storage): Not Checked

Car Oil Quantity (# of Quarts in storage): Not Checked


Water Status:

Outside Potable Water Tank Level (inches from bottom): 75% Full

Trailer Potable Water Tank Level (inches from bottom): 15 cm

Hab Potable Water Tank Level (inches from bottom): 42%

Potable Water Meter Reading: 457134

In to GreenHab Meter Reading: N/A

Notes/Comments: Green Hab not operational but water is being recycled


GreenHab:

Condition of plants in Tank 1: 5%

Condition of plants in Tank 2: Full

Condition of plants in Tank 3: 15% Full

Greenhab Notes/Comments: Green house not used


Observatory:

Telescope Used? (Yes/No): No

Observatory Notes/Comments:


Transportation:

Hab Car (NOT RENTAL): Operating

Used (yes/no): NO

Oil Checked (yes/no):

Tire Status:

Odometer at end of day:

Notes/Comments on Hab Car:


ATV 1 (Opportunity): Operating

Used: y/n YES

Oil Checked: y/n YES FULL

Fuel Consumed: Full tank then 20 minutes drive

Tire Status: Pumped Up

Comments: Driving Well


ATV 2 (Spirit):

Used: y/n YES

Oil Checked: y/n YES

Fuel Consumed: Full tank then 20 minutes drive

Tire Status: Pumped Up

Comments: Driving Well


ATV 3 (3rd Yamaha):

Used: y/n NO

Oil Checked: y/n Full

Fuel Consumed: Full

Tire Status: Needs Pumping

Comments: Battery cable connection needs repair. Purchased a connector to be installed. Installation not undertaken.


ATV 4 (Viking II):

Used: y/n NO

Oil Checked: y/n Full

Fuel Consumed: Full

Tire Status: Needs pumping

Comments:


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): A/C not operating 60 degrees F

Actual temperature upstairs (Farenheit): 58 degrees F

Thermostat setting downstairs (Farenheit): 65 degrees F

Actual temperature downstairs (Farenheit): ?


Computers/Networking Infrastructure: Operating

Notes/comments on internet/computers?: All OK operating faster compared to yesterday.

General Engineering Notes/Comments: